### v1

# Conceptualization, Data Curation Conceptualization Visualization, Writing - Review

Abstract
Condition monitoring plays a significant role in the safety and reliability of modern industrial systems. Artificial intelligence (AI) approaches are gaining attention from academia and industry as a growing subject in industrial applications and as a powerful way of identifying faults. This paper provides an overview of intelligent condition monitoring and fault detection and diagnosis methods for industrial plants with a focus on the open-source benchmark Tennessee Eastman Process (TEP). In this survey, the most popular and state-of-the-art deep learning (DL) and machine learning (ML) algorithms for industrial plant condition monitoring, fault detection, and diagnosis are summarized and the advantages and disadvantages of each algorithm are studied. Challenges like imbalanced data, unlabelled samples and how deep learning models can handle them are also covered. Finally, a comparison of the accuracies and specifications of different algorithms utilizing the Tennessee Eastman Process (TEP) is conducted. This research will be beneficial for both researchers who are new to the field and experts, as it covers the literature on condition monitoring and state-of-the-art methods alongside the challenges and possible solutions to them.
keywords:
[…]
Different deep learning and neural network-based approaches are fairly compared on a standard benchmark dataset.
Challenges regarding the data quality, explainability and possible solutions for tackling them are explored.
1 Introduction
Condition Monitoring (CM) is crucial for any industry as it improves the safety, reliability and efficiency of the systems [77]. In industrial systems, CM can be defined as the process of analyzing measurable variables to identify significant changes which are signs of a developing fault. A fault is an unpermitted deviation of the characteristic property of a variable from standard behaviour [32]. Faults can be categorized into three classes based on their time dependency, abrupt fault (stepwise), incipient fault (drift-like), and intermittent fault (happens at usually irregular intervals) shown in Figure 1. If the fault is not detected and corrected on time, it may lead to a failure which is a permanent interruption in the system performance. The tasks of fault detection and diagnosis (FDD) have various definitions in different literature; however, in this paper, we define these tasks as fault detection, fault classification, fault isolation, and fault identification, based on [32] and [33]. Fault detection is the task of determining if a fault has occurred in the system or not. Fault classification determines the type of faults, and discriminates different groups of faults. Fault diagnosis identifies the type of faults with as much specificity as possible, including the problem size, location, and time of detection. Fault isolation determines the location of each fault, for example, which part of the system has become faulty. Fault identification determines the type, magnitude and cause of the fault. In general FDD methods can be divided into three main categories, hardware redundancy, model-based fault detection, and signal-based fault detection. Hardware redundancy is a method in which multiple sensors, hardware, and software are used to measure variables. With these measurements, final FDD decisions can be made based on a weighted vote among the collected information. This method is highly reliable and common in very sensitive systems, however, it requires additional space to place the redundant hardware and can also be expensive. In model-based fault detection, faults are detected by comparing the available measurements of the system in conjunction with information from a mathematical model of the system. A residual amount is determined using the difference between actual measurements and estimates obtained from the model. Then, by applying a threshold to this residual quantity, a fault can be detected. Although this method is reliable, defining a precise model of a complex system is not an easy task, posing a challenge even to domain experts. On the other hand, signal-based methods rely solely on the analysis of data collected from sensors to detect deviations from normal behaviour and are the most frequently employed FDD methods. This approach is useful for monitoring complex systems where traditional approaches may not be sufficient, and it has been widely applied to many different industrial systems and processes. As there are numerous sensors and data in any industrial system, deep learning algorithms can be utilized to handle such information. However, this approach is highly dependent on the quality and availability of the data. The focus of this paper is on studying different state-of-the-art deep learning algorithms that have been used for the condition monitoring of industrial plants with a focus on chemical processing plants. A comparison between different approaches is established through the analysis of their application on the Tennessee Eastman Process (TEP) simulator benchmark. This paper covers different deep learning methods for condition monitoring of industrial plants and compares and summarizes these methods, also discusses the common problems of data-driven condition monitoring and introduces solutions to address them.
Condition monitoring is an inevitable task in industrial chemical systems and as such, numerous studies have been conducted in this topic. Yin et al. [104] compared basic data-driven fault diagnosis and process monitoring methods such as Principal component analysis (PCA), DPCA, two variants of Partial least squares (PLS), Fisher discriminant analysis (FDA) and Independent component analysis (ICA) on the TEP benchmark. Their study compares both the performance and the computational complexity of the examined methods. Another study by Md Nor et al. [53] provides a comprehensive review of data-driven condition monitoring methods in chemical process systems. The research covers multivariate statistical analysis such as PCA, PLS, ICA, and FDA. In addition, the work also examines machine learning approaches such as artificial neural networks (ANNs), neuro-fuzzy (NF) methods, support vector machines (SVMs), Gaussian mixture models (GMMs), K-nearest neighbours algorithm (KNN), and Bayesian networks (BNs).
Limited studies have reviewed data-driven and machine-learning FDD methods for complex industrial processes[104, 53, 105, 8]. However, to the best of our knowledge, there is no comprehensive study on deep learning and neural network-based CM approaches that fairly compares the methods based on a common benchmark case study. In this research, intelligent CM algorithms have been thoroughly analyzed and fairly compared within the context of industrial chemical processes, covering implementation challenges such as the explainability of AI methods and data scarcity. Additionally, this work considers the current state and future directions of research within the broader scope of industrial CM.
1.1 Performance Metrics
Performance and evaluation metrics serve as valuable tools for comparing and evaluating fault detection and diagnosis capabilities of different methods. This section describes the most important and widely used metrics within the context of industrial CM to provide sufficient knowledge for comparing such studies throughout the remainder of this paper. Many of these studies include the classification terms True Positive (TP), False Positive (FP), True Negative (TN), and False Negative (FN). TP represents the samples that belong to the normal class and are correctly classified as normal, FP denotes misclassifications of faulty samples as normal, TN represents the faulty samples that are correctly classified as faulty and FN refers to the incorrect classification of normal samples as faulty. It is crucial to achieve accurate detection of both normal and faulty samples. However, minimising false positives can be considered a more important task since it shows that no fault condition is mistakenly considered normal. If there is any abnormality, this prioritisation enables prompt corrective action.
Accuracy quantifies the number of correctly classified data samples out of the total data sample count, as expressed in Equation 1. It is important to note that accuracy may not be the most suitable metric when dealing with imbalanced datasets.
Precision evaluates the quality of a positive prediction made by the model and is defined by Equation 2.
Recall also known as sensitivity or true positive rate measures the number of the predicted positive instances out of all positive examples within the considered data and is expressed in Equation 3.
Fault detection rate (FDR), reflects the proportion of actual faults that are successfully detected by the system and is defined in equation 4.
False alarm rate (FAR), also known as the false positive rate indicates the proportion of falsely predicted positive instances and is presented in equation 5.
The F1-score as defined in equation 6 combines both precision and recall to provide a more balanced metric for model evaluation.
2 Research Methodology
This paper focuses on signal-based approaches for condition monitoring of industrial plants. As the amount of available sensor data can be abundant in real-world scenarios, the main focus of this research is on machine learning and deep learning techniques, especially neural networked-based approaches that can handle large amounts of data. For a fair comparison of the different applied methods, a simulated plant known as the Tennessee Eastman Process (TEP) is considered as the case study and different condition monitoring methods are compared based on their performance over this use case. A three-level keyword assembly is used to gather related papers and research to address the mentioned aspects. Table 1 shows the keyword breakdown of the query. All papers containing "Condition Monitoring", "Fault Detection", "Fault Diagnosis", "Prediction", "Fault Prognosis", and "Anomaly Detection" can be considered for review as they cover the application aspect of this paper. From the methodological point of view, the "Machine Learning" and "Deep Learning" approaches with a focus on "Neural Networks" have been selected. More than 300 papers are extracted from the Web of Science database using this query. The number of publications over the years is shown in Figure 2. Early papers on deep learning in this field were published in the early 2000s; however, these methods gained significant attention in recent years, especially after 2015, when research on more advanced machine learning and deep learning algorithms increased in popularity. This increase in popularity is largely attributed to improved computing resources, algorithmic breakthroughs, and the ability of such methods to handle complex big data. Within the scope of ML and DL, many different algorithms have been used for condition monitoring of industrial plants. Of the selected papers, the most frequently used algorithms are Autoencoders, Convolutional Neural Networks (CNNs), Recursive Neural Networks (RNNs), Deep Beleif Networks (DBNs), Attention mechanisms and Generative Adversarial Networks (GANs). Figure 3 shows the distribution of different algorithms among the studied papers.
Context |
Query |
Search Fields |
Application |
"Condition Monitoring" or "Fault detection" or "Fault diagnosis" or "Prediction" or "Fault Prognosis" or "Anomaly Detection" |
Title & Keywords & Abstract |
Method |
"Machine Learning" or "Deep Learning" or "Neural Network" |
Title & Keywords & Abstract |
Case Study |
"Tennessee Eastman" |
Title & Keywords & Abstract |
3 Tennessee Eastman Process
The Tennessee Eastman Process (TEP) is a realistic simulation of an industrial chemical plant which is the most used benchmark for process control and monitoring studies [54, 104]. The process is described in [20], and both the Fortran and Matlab simulation codes, along with numerous generated datasets, can be found online. The process consists of five main units: the reactor, the product condenser, a vapour-liquid separator, a recycle compressor and a product stripper. The flow diagram of the process is shown in Figure 4. The process produces two products from four reactants, along with an inert and a byproduct. This results in a total of eight components denoted as A, B, C, D, E, F, G, and H. The reactions are shown as follows:
The original simulation consists of 41 process variables and 12 manipulated variables, with sampling intervals of 3, 6 and 15 minutes. A total of 21 process faults, including step, random variation, slow drift and sticking faults, are defined for this simulation [104]. Refer to Table 2 for details. Some faults like 3, 9 and 15 are harder to detect.
| Fault Number | Process variable | Type | ||
| IDV(1) |
Step | ||
| IDV(2) |
Step | ||
| IDV(3) | D feed temperature | Step | ||
| IDV(4) |
Step | ||
| IDV(5) |
Step | ||
| IDV(6) | A feed loss | Step | ||
| IDV(7) |
Step | ||
| IDV(8) | A, B, and C feed composition | Random variation | ||
| IDV(9) | D feed temperature | Random variation | ||
| IDV(10) | C feed temperature | Random variation | ||
| IDV(11) |
Random variation | ||
| IDV(12) |
Random variation | ||
| IDV(13) | Reaction kinetics | Slow drift | ||
| IDV(14) | Reactor cooling water valve | Sticking | ||
| IDV(15) | Condenser cooling water valve | Sticking | ||
| IDV(16) | Unknown | Unknown | ||
| IDV(17) | Unknown | Unknown | ||
| IDV(18) | Unknown | Unknown | ||
| IDV(19) | Unknown | Unknown | ||
| IDV(20) | Unknown | Unknown | ||
| IDV(21) |
Constant position |
4 Intelligent Condition Monitoring Approaches
This section explores some of the most important algorithms that have been successfully used for the condition monitoring of industrial plants, with a focus on the TEP.
4.1 Recurrent Neural Network (RNN) and Long Short-Term Memory (LSTM)
Recurrent Neural Networks (RNN), are the networks that can learn time dependencies between data samples. These networks have feedback connections, enabling them to capture information from previous steps. The model is usually trained with back-propagation and suffers from gradient vanishing/exploding. Long Short-Term Memory (LSTM) networks are improved RNNs, consisting of memory cells and the gating mechanism for capturing information over a long time. The forget gate handles the gradient vanishing problem [29]. Gated recurrent unit (GRU) solves the vanishing gradient problem like LSTM but with fewer gates and no separate internal memory. RNN, LSTM and GRU structures are shown in Figure 5.
One of the first researches that used LSTM for chemical process fault diagnosis is [126], in this case, no specific feature extraction is needed. Applying batch normalization makes the convergence of the network much faster. To improve the diagnosis using LSTM networks, Han et al. [27] used an iterative method to optimize the number of hidden layer nodes. Kang et al. [36], used RNNs for fault classification. It is shown that RNN can create an even spatial distribution of fault types and classify them accurately, for TEP data a 4-layer RNN could reach the average classification accuracy of 92.5% for all faults, however, it can not detect fault 15. To overcome this problem LSTM network is used and the average accuracy increased to 95%. Compared to the RNN, LSTM has an additional memory cell and three extra gates. To extract long-term dependencies in raw data and detect abnormal values, a Stacked Long Short-Term Memory (LSTM) network with three layers LSTM, is introduced in [118]. As the historical information is stored in the memory, the monitoring system can operate in real time. The average FDR for all TEP faults is 99.49%. Chadha et al. [13] used bidirectional recurrent neural networks for fault diagnosis, enabling the addition of more information about adjacent points into a training sequence than their unidirectional counterparts. They introduced a novel data preprocessing and restructuring for efficient network training. The average FDR for 21 faults of TEP is 88.14%. To overcome nonlinearity, and, the high degree of spatio-temporal correlations, and incomplete characterization of the uncertainty in process noise and dynamics, a novel Bayesian recurrent neural network (BRNN) BRNN-based FDI method is proposed in [80], where we can have a probabilistic model of input-output mapping. This method can be used for fault propagation path analysis as well. To improve the feature extraction of unidirectional RNNs that only consider positive direction, bidirectional RNN (BiRNN) models are used by [120] for FDD of TEP. Simple RNN unit (SRU), LSTM, and GRU are used in making bidirectional networks with the average FDR of 91.0%, 92.3% and 92.7% respectively.
Mirzaei et al. [56], compared LSTM and GRU networks for FDD of chemical processes by using t-SNE visualization. GRU performed better than LSMT, especially for fault 15 which is very hard to detect. The average accuracy of LSTM and GRU for all faults are 95% and 95.9% respectively.
Yuan et al. [113], used GRU to extract the information from correlated sequential data, a moving horizon is used at the input of the GRU to divide the raw data into the sequence units. To overcome the covariate shift and improve the network’s learning efficiency, batch normalization is used. BN-based GRU is faster and more accurate. GRU and BN-GRU accuracy is 83.42%, and 87.36% respectively.
4.2 Convolutional Neural Network (CNN)
Convolutional Neural Networks (CNNs) were introduced in the late 1980s and initially used to detect patterns in images [37]. These powerful feature extraction networks are widely used for fault detection. A schematic of a CNN is shown in Figure 6.
One of the first applications of using CNN for TEP fault detection and diagnosis is introduced in [92]. In this approach, a deep convolutional neural network (DCNN) with dropout extracts features from time series data. The implemented method reached the average FDR of 88.2% for 20 faults of TEP, however, this model highly relies on historical fault data samples and is not very effective for super complex chemical processes with thousands of variables. Some modifications to the CNN network are done in [12] to increase the generalization ability of the model, improving feature extraction abilities and reducing the overfitting such as residual connections, inception module, and dilation convolution. With these improvements, the F1 score of hard detectable faults 3,9,13,15 is improved to 0.91 . Frequency domain data contain valuable information that is usually neglected in the traditional FDD of chemical processes. In [116], the Fast Fourier Transform (FFT) is used for TEP feature extraction, and the output data is fed into a convolutional neural network which is used as a binary classifier. The FDR of this model on TEP data considering all faults except 3,9,15,16 is 96.9%. In CNN all the features have the same importance in the classification process. Multi-head attention CNN (MACNN) [18] are introduced to distinguish the importance of different features, where the extracted features from convolution layers are fed into the multi-head attention (MA) layers, finally a softmax classifier is used to classify the different faults, the performance is improved compared to feedforward neural network (FNN) and CNN models. Multiblock temporal convolutional network (MBTCN) for FDD is introduced in [28] using the idea of "local extraction and global integration", it considers the cross-correlation and the temporal-correlation in the multivariate processes’ data. In this method, most related process variables are divided into the same sub-blocks where TCN extract the features. It achieved 96.2% accuracy in TEP considering all fault classes but 3,9,15. In [111], a multichannel one-dimensional convolutional neural network (MC1-DCNN) was used for FDD of the TEP process using extracted features from the Wavelet transform. Wavelet transform is a powerful tool for extracting time and frequency domain features. The average recognition accuracy of the implemented method for all 21 faults of TEP is 94.08%. To handle multi-time scale data, Gao et al. [21], introduced a feature extraction method based on CNN. The similarity measure is used to eliminate fault-irrelevant variables. The fixed multiple sampling (FMS) method is introduced for converting a feature map with an arbitrary shape into a fixed-length vector to handle different time scales of data. The average FDR for all 21 classes using the multi-time scale model is 98.53%, while the single-time scale model is 94.65%. A multi-model monitoring method is introduced in[39], in which first a CNN model detects some faults quickly and with minimum computational power, then a wavelet-assisted secondary CNN model accurately detects the remaining faults and also filters the noise and redundant information, the FDR of the model for 20 faults of TEP is 93%. A normalized CNN is used in [86] to overcome the problem of gradient explosion and disappearance, moreover the convergence speed of the network is improved. This network consists of the depth-normalized convolution network and improved second-order pooling for feature fusion. The average fault detection accuracy for all faults is 92.75%. It should be noted that this method needs a large amount of calculations. To deal with high dimensional nonlinearity the combination of matrix diagram and multi-scale convolutional neural network (MsCNN) is used in [79]. Time series data is converted into matrix diagrams (2-D images). The average fault detection accuracy for all faults is 88.54%. It should be noted that running this model is very time-consuming. Sparse one-dimensional convolutional neural network (S1-DCNN) [112], is introduced to extract features from complex signals, the method uses a sparsity regularization layer at the output of the convolution network to remove any redundant representations and extract fault-discriminative features. In [71], adaptive multiscale convolutional neural network (AMCNN), two attention modules, adaptive channel attention and adaptive spatial attention, are incorporated to select the most useful features and suppress redundant information. Another novelty in this paper is the proposed optimization method using a weighted loss function consisting of classification loss, regularization loss and triplet loss to improve robustness and generalization. the average fault detection rate for 20 faults of TEP is 93.02%. Dynamic convolutional neural networks can extract temporal correlation of process data and are useful for feature selection, and feature sequence arrangement as in [19], first, the genetic algorithm (GA) is used for feature selection and then CNN is used for fault diagnosis. The FDR for 20 faults of TEP is 89.72%. A 1D-CNN network is used in [58] for FDD. This model exploits the correlation in the data, and the average FDD accuracy for all TEP faults is 93%. Multiblock adaptive convolution kernel neural network (MBCKN) [26] is a feature extraction method for large-scale processes, in which the process is decomposed into several blocks, and the attention mechanism is used for convolutional kernel selection for each block. Finally, an adaptive fusion method fuses the features of different blocks. This model can detect 21 faults of the TEP with an average accuracy of 94.15%.
4.2.1 LSTM and CNN
Yuan et al. [114] proposed a framework for feature extraction and fault diagnosis, where the discrete wavelet transform and CNN are used for feature extraction then the LSTM network reduces the useless information and finally classification is done by using a softmax classifier. The model is validated on a few faults (Fault 1, 2, 4, 8, 10, 12, 13, 14, 17, and 18) of the TEP and the average F1-score is 0.9873. In [76], first, the LSTM network is used for feature extraction, then convolutional kernels extract characteristics of the data. This combination results in the high average fault detection accuracy of 92.06% for the TEP data set. This model can detect faults 3, 9, and 15 which are very difficult to diagnose with a very high accuracy. Bao et al. [6] used a deep convolutional neural network (DCNN) to extract spatial features from raw data and then fuse features into the bidirectional recurrent neural network (BRNN). No additional data preprocessing is needed in this case. The average accuracy for 20 faults of TEP is 91.7%. The combination of the sliding window on integrated 2D image-like data of feature information and time delay information, and the CNN-LSTM model is used in [31] for process fault detection. This method is robust against noise and converges faster compared to separate CNN and LSTM networks.
4.3 Auto-encoders
Autoencoders are models first introduced in the late 1980s. They consist of two networks, the encoder and the decoder where the encoder maps the input data to a hidden layer and the decoder tries to reconstruct the input data, in general, the process is done unsupervised [75]. They are powerful tools for feature extraction and dimension reduction. The output of the encoder is the feature representation and can be used for further analysis like as the input of a classifier. Different types of autoencoders are shown in Figure 7.
Autoencoders have been widely used in condition monitoring and fault detection and diagnosis of industrial plants where the available data is massive, specifically for chemical processes and TEP. One of the first attempts of using autoencoders for fault diagnosis of the TEP is done in [49], a stacked sparse autoencoder (SSAE) is implemented for extracting and learning high-level features and underlying fault patterns which is very useful for detecting incipient faults, moreover time correlations among the samples is also considered, unlike other methods that have poor performance in diagnosing incipient faults, this algorithm could classify faults 3,9,15,21 of the TE which are very hard to detect with the accuracy of 95.75%, 99.25%, 98.75%, 99.5% respectively. Another powerful tool for fault detection especially incipient faults in industrial plants is the Deep LSTM Supervised Autoencoder Neural Network (Deep LSTM-SAE NN) [1]. The assumption that having different models for different faults increases the sensitivity of the model is made in the implementation of this method. The average accuracy for all the classes is 93.23%.
Stacked denoising autoencoder (SDAE) alongside the k-nearest neighbour (KNN) rule is used for automated feature learning in [124], the SDAE automatically extracts crucial features from nonlinear process data. Neural component analysis (NCA) for fault detection is introduced in [125], which is a unified model including a nonlinear encoder and a linear decoder, the network is trained with orthogonal constraints such as those used in PCA to alleviate the overfitting problem in autoencoders. In [15] a variational recurrent autoencoder is used for process fault detection. This method monitors the process in probability space, to handle processes nonlinearity more efficiently. Recurrent neural networks are also used to consider temporal dependency between variables. The negative variational score (NVS) is introduced as the monitoring metric, which considers reconstruction error and similarity between prior and posterior probability distributions. The model is validated on the TEP and the average fault detection rate of all faults expect 3,9,15 is 96.3%. The combination of an autoencoder and LSTM is a powerful tool for fault detection of rare events. In [61], an autoencoder is trained with offline normal data and used as an anomaly detector. The predicted faults will be classified by using an LSTM network. The average fault detection accuracy for the first 20 faults is 91.9%, and by not considering 9 and 15 faults which are very hard to detect, the average accuracy is 96.8%. The time delay of the fault detection of each class is also studied in this work. In [103], a fault detection model based on Mutual Information and Dynamic Stacked Sparse Autoencoder (SSAE) is introduced. The variables with a strong autocorrelation and their lag time of maximum autocorrelation were chosen using time series input data. Then, dynamic augmented data is fed to the SSAE for feature extraction and learning latent information in the dynamic process data. Based on selected process variables the average accuracy of fault detection of all 21 faults is 85.06%. Liu et al. [43], introduced a denoising sparse autoencoder (DSAE) to address the noise interference and nonlinearity in process variables. This method can achieve an accuracy higher than 99% for some faults. There are usually multiple operating modes in a chemical process, and fault detection models that are built based on one mode may have high false alarm rates. For multimode process monitoring and handling unseen modes a self-adaptive deep learning method based on local adaptive standardization and variational auto-encoder bidirectional long short-term memory (LAS-VB) is introduced in[93], in this research unstable tendency in the local moving window is detected. This method is not efficient for very slow-drif faults. Combining autoencoders and clustering algorithms for fault detection is a powerful tool, especially for dealing with unlabeled data. Zheng et al. [127], used a convolutional stacked autoencoder (SAE) network for feature extraction. Feature visualization is done by the t-distributed stochastic neighbour embedding (t-SNE) algorithm and DBSCAN clustering. The implemented method could detect Faults 1, 2, 4, 6, 7, 11, 13, 14, 17, 19 and 20 with average accuracy of 93%. Clustering is done unsupervised, however, a knowledge-based cluster annotation is used for labelling. Chadha et al. [11] also used a convolutional autoencoder for handling unlabeled data. The latent features are split into discriminative and reconstructive, and an auxiliary loss based on k-means clustering is used for discriminatory latent variables. The classification average accuracy for 3, 9, 15, and 21 faults which are very hard to detect is 50.04%. In [4], a source-aware autoencoder (SAAE) for detecting unseen faults and handling imbalanced data sets is introduced that uses Bidirectional long short-term memory (BiLSTM) and residual neural network (ResNet) to deal with randomness. The training can be done by using just normal data, and then fine-tuning by using emerged faulty samples. The proposed method detects unseen faults with an accuracy of 93.8%, moreover, the average accuracy of fault detection for different fault-to-normal ratios is 91.8%. In the balanced condition, the average accuracy is 94.3%. As usually there are not enough labelled samples, semi-supervised learning is developed that uses unlabeled data to help with the limited labelled data during model training. A consistency regularization autoencoder (CRAE) framework is introduced in [51], which captures correlations from labelled and unlabeled data. Previous samples are augmented with the current samples to increase the fault detection speed and accuracy. By just having the labels of just 20% of the dataset, the average fault detection accuracy of all the faults in the TE dataset except 3, 9, and 15 is 92.95%.
Peng et al. [66], introduced a non-revisiting genetic cost-sensitive sparse autoencoder(NrGCS-SAE), for solving the imbalanced fault diagnosis problem. The genetic algorithm optimizes class weights while the non-revisiting strategy increases exploration ability. This method could detect faults in the unbalanced TE dataset with an average accuracy of 92.02% while the normal to-fault ratio was 10:1. Convolutional autoencoders are powerful feature extraction tools, however, the pertaining phase is done unsupervised, to improve the extracted features and make sure they contain internal spatial information, supervised convolutional autoencoder (SCAE) is introduced [106]. The average accuracy of this method on the TE dataset is 96.19%, and it could reach an accuracy of about 95% in only two epochs. One-dimension residual convolutional auto-encoder (1DRCAE) is a model introduced in [107], for unsupervised feature extraction, where two residual learning blocks are embedded between the encoder and decoder to improve the speed and accuracy. The fault detection rate of the model on the TEP dataset on all faults except 3, 9, and 15 is 90.8%. To utilize unlabeled data more efficiently, Zhang et al. [122], combined LSTM for temporal feature extraction and ladder autoencoder (LAE) for semi-supervised learning. LSTM-LAE achieves the interpretability to extract fault-relevant process variables. By just having the labels of 11% of the dataset, this model achieved the average accuracy of 95.9% for all faults except 3,13, 15 and 16. Stacked spare-denoising autoencoder (SSDAE)-Softmax[44], is introduced for robust fault detection in complex industrial processes. The state transition algorithm (STA) is a global optimization method that is used for hyperparameter optimization. The fault detection accuracy of the model on the TEP reaches the average of 96.6% for faults 1, 2, 4, 5, 7, 13, and 17.
Conventional autoencoder algorithms usually focus on the reconstruction error and neglect lots of useful information in the latent space for FDD. To address this problem one-dimension convolutional adversarial autoencoder (1DAAE) is introduced in [85], which uses the adversarial thought for clustering the latent variables to a prior distribution. Both reconstruction error and latent variable distribution are considered in the training of the model. The mean accuracy for all 21 faults of TEP is 86%. To monitor the nonlinear process and capture the process dynamics, a novel dynamic-inner convolutional autoencoder (DiCAE) is introduced in [121], that integrates a vector autoregressive model into a 1-dimensional convolutional autoencoder. The average FDR of 96.8% is achieved on the TEP for all the classes except 3, 9 and 15. Jang et al. [34], combined variational autoencoder and a generative adversarial network proposing an adversarial autoencoder (AAE), to generate features that follow the designed prior distribution. These features improve the reliability and stability of FDD algorithms. The average FDR for all 21 faults of TEP is 78.55%. This method decreases the fault detection delay and can detect faults faster compared to VAE. To reduce the irrelevant information with the faults during the feature extraction, a neuron-grouped stacked autoencoder (CG-SAE) is introduced in [60] hierarchical fault-related feature representation. In this method, hidden neurons are divided into different groups representing the data categories, to extract category-related features. This method achieved the average FDR of 96.02% for all faults except 3, 9 and 15. Li et al. [41], introduced a multimode feature fusion algorithm that selects the most relevant variables for each fault using the minimum redundancy-maximum relevance (mRMR) method, and then uses a stacked autoencoder for feature extraction. The average FDR for all faults except 3,9,15 is 88.04%, which is improved compared to SAE with the FDR of 76.77%. To reduce the noise in high-dimensional process signals, a one-dimensional convolutional auto-encoder (1D-CAE) is introduced in [14], This model does the feature extraction in a hierarchical structure, and this model outperforms SVM, DBN and SDAE. Jin et al. [35], used neurons with wavelet activation function (wavelons) and exponential linear units (ELU), to make a wavelons-constructed autoencoder-based deep neural network (WA-DNN), the average FDR for all 21 faults is 95.86%.
Traditional fault detection techniques do not take quality indicators into account, they only look at process data. A mixing stacked auto-encoder, consisting of a nonlinear encoder and a linear decoder, is introduced in [98] to extract features of process variables and quality indicators while keeping feature extraction and model construction separate. Prior knowledge assists in enhancing the predictive accuracy of quality indicators and makes models easier to interpret in order to obtain more dependable monitoring performance. Stacked denoising auto-encoders (SDAE), are used to extract nonlinear latent variables and features from noisy data, however, the dynamic relationships are neglected. To consider them, a recursive stacked denoising auto-encoder (RSDAE) is proposed in [130]. For extracting more effective features from time domain process signals, a convolutional long short-term memory autoencoder (CLSTM-AE) is introduced in [108], where a residual block is used to select key information and improve the accuracy and reduce gradient disappearance and gradient explosion in the training phase. FDR for all the faults of TEP is 91%. Nonlocal and global preserving stacked autoencoder (NLGPSAE) [101], can extract structure-related features from nonlinear process data by using a regularized objective function that can control the reconstructed space. The average FDR for all the faults of TEP considering and SPE statistics are 94.8% and 91.8% respectively.
To consider geometrical information from process signals during the feature extraction of SAEs in fault detection, manifold regularized stacked autoencoders (MRSAE) are introduced in [110] that considers the neighbourhood information. MRSAE does not need a large amount of labelled data. This model is superior in detecting certain types of faults like 5, 16, and 19, and the average FDR for 21 faults of TEP is 91.42%. To achieve interpretable analysis along the fault detection task, a novel orthogonal self-attentive variational autoencoder (OSAVA) model is introduced in [10]. Orthogonal attention extracts the correlations between variables and temporal dependency of different timesteps. The self-attention mechanisms utilized in the VSAE consider information from all timesteps to improve the performance. This model has a low detection delay an average of 24.2 minutes, the average FDR is 94.7%. To achieve effective and interpretable features for condition monitoring, a slow feature analysis-aided autoencoder (SFA-AE) is introduced in [38], which extracts deep slow variation patterns from the extracted features of AE. These slow patterns can show abnormal states. The approach uses a neural network consisting of a process variable selection (PVS) block, followed by an encoder-decoder structure and a deep slow feature analysis (DSFA) section. The PVS block utilizes 1D convolutional layers to generate an attention map that represents feature importance for selecting crucial process variables. Next, the encoder and decoder senctions use convolutional LSTM and deconvolutional LSTM layers respectively to learn spatiotemporal relationships of the input variables. The average FDR for 21 faults of TEP is 94.87%.
4.4 Generative Adversarial Network (GAN)
Generative Adversarial Networks (GAN) are deep learning models utilized for dealing with imbalanced data as they are generative algorithms that can increase the number of samples and improve data-driven algorithm accuracy by providing more data. GANs consist of two networks, a generator model that tries to generate synthetic samples and a discriminator model that discovers if a sample is generated by the generator or not[22]. A schematic of a GAN is shown in 8.
For solving the imbalance fault diagnosis problem an enhanced auxiliary classifier generative adversarial network (EACGAN) is introduced [65]. In this model, the boundary-seeking loss is used for stabilizing the training process, with the normal to-fault ratio of 10:1 and considering all faults in the TEP dataset except 3, 9 and 15, the accuracy of the model is 77.01%. High-efficiency GAN model (HGAN) [72] is a model that integrates Wasserstein GAN and Auxiliary Classifier GAN to improve the stability and efficiency of the training process. This model is not just limited to generating data but directly can be used for fault classification. For the balanced TEP dataset, considering all faults the accuracy is 86.6%, while by the normal to fault samples ratio of 5:1, the accuracy is 82.0%. Sparse semi-supervised GAN (SSGAN) [45] is a model that uses unlabeled and labelled data for training the model. In this model, the discriminator is changed into a multi-dimensional classifier and the loss function for each labelled, unlabeled and generated type of data is defined separately, and the total loss is used in optimization. Lomov et al. [46], studied different recurrent and convolutional networks and introduced a novel temporal CNN1D2D architecture. A GAN is used to enrich the training data samples. This method performs well in the detection of 3, 9, and 15 faults, which are very challenging.
4.5 Deep Belief Network (DBN)
Deep Belief Networks are primarily used for unsupervised learning tasks, especially for feature learning, they are made of multilayer Restricted Boltzmann machines (RBMs), which are energy-based probabilistic models. DBN consist of visible and hidden layers, the visible layer represents the input data, while the hidden layers capture abstract and complex features. Each layer is fully connected to the adjacent layers, but there are no connections within a single layer. A schematic of DBN is shown in 9 .
One of the first researches that used DBN for chemical plants FDD is [95], unlabeled data is used to pre-trian the DBN, and then labelled data is used for training the DNN using backpropagation. The average classification accuracy is 80.5%. Using deep learning techniques for layer-by-layer feature extraction may lead to the weakening or vanishing of the fluctuations that reflect the fault information. To address this issue, in [109], the support vector data description (SVDD) and moving average filter technologies are utilized, moreover, in the structure of DBN, the traditional Restricted Boltzmann Machine (RBM) is replaced with the Gaussian Restricted Boltzmann Machine (GRBM). To reduce overfitting and improve the generalization of DBN, dropout is used in DBN [89]. Dropout randomly disables a certain portion of units in the training process. By using this method the average F1-score for all faults of TEP is 0.81. By using deep learning for FDD, important information in the raw data might be filtered during feature compression, to solve this problem an extended deep belief network (EDBN) is proposed in [88], where each extended restricted Boltzmann machine (ERBM) receives raw data and hidden features as inputs during the pre-training phase. The dynamic properties of process data are then taken into account while building a dynamic EDBN-based fault classifier. The average FDR of all TEP faults except 3,9,15 using EDBN is 94.31%, which is slightly better than DBN 93.89%. The existing noise level in industrial environments affects the performance of deep learning FDD models. To reduce the noise effects, a penalty factor is added to the network structure, avoiding the local optimal situation of a DBN and increasing FDD accuracy, besides an adaptive lifting wavelet (ALW) that reduces the noise level leading to ALW-DBN model[102]. The average FDR of TEP for all faults is 97.33%. However, The ALW-DBN model’s training takes longer because of the iterative penalty optimisation algorithm. Additionally, the DBN method has obvious drawbacks when attempting to extract the time information from chemical data in comparison to the DCNN method. Lui et al. [42], introduced an index for measuring fault information of high dimensional features and finding deep highly-sensitive features (DHSFs) from the extracted features of a DBN. In the mentioned research Gauss– Bernoulli Restricted Boltzmann Machine (GRBM) is used to replace the traditional binary Restricted Boltzmann Machine (RBM), to reduce information loss. Finally, Euclidean distance (ED) is used to calculate the distance between normal and fault samples and the moving average (MA) to eliminate the effect of noise. The average FDR of the proposed DHSF-DBN method is 88.67%. Yang et al. [100] proposed a pairwise graph regularized deep belief network (PG-DBN) model for FDD. Two different graph constraints are considered on the hidden layer of the RBM, preserving the feature manifold structure in the same class of the data and penalizing the feature manifold structure in different classes of the data, respectively. The average classification accuracy of PG-DBN, G-DBN, CNN and DBN which are studied in this paper is 93.87%, 91.98%, 89.18% and 83.1% for all 21 faults of TEP.
4.6 Transformers and Self-attention Networks
Transformer models were introduced by Vaswani et al. [83] in 2017 and are the foundation for many state-of-the-art NLP models. They consist of an encoder and a decoder, and the input and output are sequences. The innovation of this model is in the attention mechanism, which weights the importance of different samples in the input sequence, so the model can capture important information from the entire input sequence rather than a fixed window size. The transformer structure is shown in Figure 10.
Fault happens when a variable deviates from the normal condition. Rapid changes can easily be captured by ML models, however, slow changes require a long time series, that can be analysed by attention-based models that have a global view of the whole process. Deep learning models like RNNs and CNNs have been widely applied for FDD of chemical processes, however, they are insufficient in extracting features of long-term dependencies. Self-attention mechanisms and transformer models can perform better for this goal. These models are primarily designed for natural language processing (NLP) tasks, however, time series data are very similar to text sequences [90]. One of the first researches that used transformers for chemical process fault detection is [117], they introduced a general transformer in which self-attention is replaced by graph attention, and it achieved an accuracy of 80.10% for 21 fault conditions. Wei et al. [90] proposed a Target Transformer algorithm, that has both self-attention and target-attention mechanisms. The model retains the original transformer’s encoder and applies target attention to the decoder to extract fault-related features. The average FDR for 20 faults of TEP is 90.39%. Zhou et al. [129] proposed an Industrial Process Optimization ViT (IPO-ViT) model based on Vision Transformer (ViT) for FDD. In this research, CNN which can extract local features and information is combined with multi-head attention which is famous for its global attention. FDR of the proposed model is increased compared to CNN and LSTM models, especially for faults 3,9 and 15. The average FDR for 20 faults of TEP is 94.08%.
Zhang et al. [119], applied the self-attention (SA) mechanism to the gated recurrent unit (GRU), and introduced an encoder-decoder framework for FDD, where the encoder extracts the features and the decoder does the fault detection. The model is validated on TEP and the average recall rate for GRU-SA is 0.9504 which is better than GRU’s 0.9214. However, this approach can only be used when enough and reach fault samples are available. Xiong et al. [96], introduced a deep learning fault diagnosis method based on fully convolutional networks, long-short-term memory, and attention mechanism called attention-based LSTM and Fully Convolutional Network (ALSTM-FCN) for FDD. The CNN and LSTM layers are incorporated for their spatial and temporal feature extraction abilities respectively, and the attention mechanism is included for its ability to adaptively focus on the influence of important features. Dropout is used to prevent overfitting, and batch normalization to avoid the negative effects of covariate shift. Results show that the proposed method performs better in comparison to traditional CNN and LSTM models. The average accuracy for all 20 faults of TEP is 88.80% and for faults 3 and 9 less than 80%.
A summary of various algorithms that were successfully used for condition monitoring of TEP, along with their specifications is presented in Table 3.
| Model | Reference | Citation | Specification | Faults Information | Average Accuracy |
LSTM |
[114] |
Yuan 2019 |
Discrete wavelet transform and CNN for feature extraction and softmax classifier |
1, 2, 4, 8, 10, 12, 13, 14, 17, and 18 |
98.73% |
BiRNN |
[120] |
Zhang 2019 |
Bidirectional recurrent neural networks using gated recurrent unit (GRU) |
All |
92.70% |
LSTM + CNN |
[76] |
Shao 2019 |
First LSTM and then CNN are used for feature extraction. The detection of faults 3, 9, and 15 are improved. |
All |
92.06% |
RNN |
[36] |
Kang 2020 |
Using a 4 layer RNN, however, could not detect fault 15 |
All |
92.50% |
LSTM |
[36] |
Kang 2020 |
Better performance compared to RNN speciallly for fault 15 |
All |
95% |
LSTM |
[118] |
Zhang 2020 |
Three layers LSTM are stacked |
All |
99.49% |
BiRNN |
[13] |
Chadha 2020 |
Bidirectional recurrent neural networks |
All |
88.14% |
GRU |
[56] |
Mirzaei 2022 |
The detetcion of fault 15 is enhanced compared to LSTM |
All |
95.90% |
CNN |
[92] |
Wu 2018 |
CNN with dropout |
20 faults |
88.20% |
CNN |
[116] |
Zhang 2019 |
FFT for feature extraction |
All faults except 3,9,15,16 |
96.90% |
CNN |
[21] |
Gao 2020 |
Handling multi time scale data using fixed multiple sampling method |
All |
98.53% |
CNN |
[28] |
He 2021 |
Multiblock temporal convolutional network (MBTCN), dividing process variables into subblocks |
All faults except 3,9,15 |
96.20% |
CNN |
[111] |
Yu 2021 |
1D CNN and Wavelet transform for feature extraction |
All |
94.08% |
CNN |
[86] |
Wang 2021 |
Normalized CNN to improve the stability and speed of the training |
All |
92.75% |
CNN |
[26] |
Guo 2022 |
Multiblock adaptive convolution kernel neural network (MBCKN) for large scale process using attention mechanism |
All |
94.15% |
Autoencoder |
[35] |
Jin 2018 |
Using wavelet activation function in the autoencoder |
All |
95.86% |
Autoencoder |
[108] |
Yu 2020 |
Convolutional LSTM memory autoencoder (CLSTM-AE) and residual block for increasing the accuracy and make the training process more robust |
All |
91% |
Autoencoder |
[110] |
Yu 2020 |
Manifold regularized stacked autoencoders (MRSAE) considering neighbourhood information needs a few labelled data |
All |
91.42% |
Autoencoder |
[121] |
Zhang 2022 |
Vector autoregressive model and 1-dimensional convolutional autoencoder for capturing the process dynamics |
All faults except 3, 9 and 15 |
96.80% |
Autoencoder |
[60] |
Pan 2021 |
Neuron-grouped stacked autoencoder (CG-SAE) as a hierarchical feature extractor for different groups of faults |
All faults except3, 9 and 15 |
96.02% |
Autoencoder |
[10] |
Bi 2021 |
Orthogonal self-attentive variational autoencoder (OSAVA) to increase the interpretability |
All faults except 15, 16 |
94.70% |
Autoencoder |
[38] |
Li 2021 |
Slow feature analysis-aided autoencoder (SFA-AE) to deal with slow variations in the data |
All faults except15, 17 |
94.87% |
Autoencoder |
[1] |
Agarwal 2022 |
Combining LSTM and Supervised Autoencoder |
All |
93.23% |
Autoencoder |
[15] |
Cheng 2019 |
Variational recurrent autoencoder combines RNN and VAE |
All faults except 3,9,15 |
96.30% |
Autoencoder |
[106] |
Yu 2022 |
Supervised convolutional autoencoder (SCAE) for better feature extraction |
All faults except 3,9,15,16 |
96.19% |
Autoencoder |
[85] |
Wang 2023 |
Convolutional adversarial autoencoder (1DAAE) |
All |
86% |
Autoencoder |
[101] |
Yang 2023 |
Nonlocal and global preserving stacked autoencoder (NLGPSAE) using a regularized objective function |
All |
94.80% |
DBN |
[88] |
Wang 2020 |
Extended deep belief network (EDBN)enhances the data quality during feature compression |
All faults except3,9,15 |
94.31% |
DBN |
[102] |
Yao 2021 |
Adaptive lifting wavelet-DBN (ALW-DBN) to reduce the noise level |
All |
97.33% |
DBN |
[100] |
Yang 2021 |
Pairwise graph regularized deep belief network (PG-DBN) to monitor and control the feature manifold structure of the data |
All |
93.87% |
Transformer |
[117] |
Zhang 2022 |
Self attention transformer |
All |
80.10% |
Transformer |
[90] |
Wei 2022 |
Target transformer using both self-attention and target-attention mechanism |
20 faults |
90.39% |
Transformer |
[129] |
Zhou 2023 |
Vision Transformer and CNN Increasing the accuracy, especially for faults 3,9 and 15 |
20 faults |
94.08% |
Attention mechanism |
[119] |
Zhang 2019 |
Combines self-attention (SA) mechanism to the gated recurrent unit (GRU) |
20 faults |
95.04% |
Attention mechanism |
[96] |
Xiong 2023 |
Attention-based LSTM and Fully Convolutional Network (ALSTM-FCN) with Dropout |
21 faults |
88.80% |
| Method | Advantages | Limitations |
Recurrent Networks |
Process sequential or temporal data, modelling time dependencies |
Prone to gradient vanishing/exploding |
Attention Models |
Handling sequential data and weighting important samples in the input sequence. No fixed window size is needed. Can capture slow changes in long time series and extract features of long-term dependencies |
Needs rich historical data |
CNN |
Powerful feature extraction capability. Has denoising capabilities [47]. |
May need multiple layers and large labelled datasets to perform well |
Autoencoder |
Feature extraction and dimension reduction ability. Unsupervised training. Robust against noise |
Pretraining needed |
GAN |
Generate data to enhance the data imbalance problem. Can be used for semi-supervised classification [72] |
Stability problem in the training |
DBN |
Can be trained unsupervised or supervised. Energy-based probabilistic models |
Computationally expensive training |
5 Discussion, Challenges and New Frontiers
This paper provides a summary of the most important deep learning algorithms that have been used for condition monitoring of industrial plants. Deep learning is a powerful tool for feature extraction and handling the large and complex data of chemical plants. LSTM and CNN models can find the hidden representations of the data. In general recurrent networks such as RNN, LSTM and GRU perform well on time series, sequential data and data with time dependencies. Transformers and attention mechanisms are also tools for analysing sequences by weighting the most important samples. Autoencoders reduce the dimension of the data and extract features. Different architectures and models, such as CNNs, can be used to design the encoder and decoder parts. Autoencoders are unsupervised learning methods and some of the variations like denoising Autoencoders are robust against noise. GANs are designed to generate data and can be used to enrich the dataset and increase the accuracy of the condition monitoring. A summary of the discussed methods and their advantages and limitations is provided in Table 4. The following part of this paper discusses challenges of data-driven condition monitoring. Some challenges like unlabeled, unseen or imbalanced data are related to the availability of the data, which can be handled by choosing the proper ML algorithm. Using AI and ML algorithms is consistently challenging due to their black-box nature. The system’s decisions are not easily interpretable for humans, posing difficulties in establishing trust and hindering real-world implementation. To address this problem, explainable AI, causal AI and ensemble learning approaches are introduced as a way to increase the accuracy and trust of the decision-making systems.
5.1 Unlabeled Data
Collecting and labeling the data is challenging, especially in many industrial setups where even if the data is collected properly, labeling it may not be straightforward. Labelling the data is a time-consuming and expensive task that usually needs the supervision of an expert. Different approaches like unsupervised learning and Self-supervised learning have been used to deal with this problem. If only the normal data is labelled, anomaly detection methods can be useful.
Most FDD methods fail when there is not enough labelled data. Peng et al. [67], proposed a method that uses a novel form of the bidirectional gated recurrent unit (BGRU), weighting each training example to overcome the class imbalance and cost-sensitive active learning for selecting the candidate instances to explore the unlabelled data. The proposed method is validated on TEP, by an imbalance ratio of 0.5 and the unlabelled proportion data of 0.6 and achieved the average FDR of 97.5%. Deep learning methods usually need large labelled datasets to achieve good performance. Self-supervised learning is an approach that can learn features from unlabeled data. Pöppelbaum et al. [70], proposed a novel self-supervised data augmentation technique based on SimCLR contrastive learning. The goal of contrastive learning is to compare different samples without any labels, the data itself provides supervision. In the SimCLR framework, multiple data augmentation methods for feature representation are used. By only considering 5% of the available training data, the classification accuracy is 80.80% for all TEP faults. To address situations where there are only a few labeled samples of normal data and the remaining available data, including abnormal samples, is unlabeled (positive-unlabeled (PU) learning), a common problem in the industry, a three-step high-fidelity PU (THPU) approach for FDD is introduced in [128]. The approach utilizes self-training nearest neighbors (STNN) to generate a positive dataset from unlabeled training data. Subsequently, a positive and negative data recognition (PNDR) algorithm, based on the SAE, augments the positive dataset and generates a negative dataset based on the unlabeled data. Finally, CNN is used as a binary classifier. With only one labelled normal sample, the FDR of the proposed method on TEP is 92.32% and will increase to 93.75% with 50 labelled normal samples. As Faults 3, 5, 9, 15 and 16 are very similar to the normal operational state, they were not considered in this study. A semi-supervised anomaly detection method composed of an autoencoder and LSTM model is introduced in [57], in this method, a control chart called Open Up is used for fault diagnosis. Open Up provides visual support that enables the interpretation and diagnosis of detected faults in a production system. The average accuracy of the model on a subset of TEP with 6.25% labelled data is 97.46%. Gorman et al. [23], proposed an interpretable anomaly detection method based on 1-Dimensional Convolutional Autoencoders (1dCAE) and Localised Reconstruction Error (LRE). In LRE each input channel is evaluated separately, and the sensors and data that lead to anomalies are identified, outperforming global reconstruction errors.
Self-supervised contrastive learning extracts information from a large amount of unlabeled data. To enhance the performance of fault diagnosis in scenarios with numerous normal samples and only a few fault classes, a long-tailed distribution is employed in [63]. Progressively balanced supervised contrastive learning (PBS-SCL) is utilized to generate a balanced data batch, facilitating feature learning. Finally, a learnable linear classifier (LLC) is used for classification. The model was validated on a subset of TEP data. By having just 1,5,10,15, 20 samples of faults for each category and enough normal samples, the classification accuracy is 47.72%, 63.92%, 72.23%, 86.90%, and 86.59% respectively.
5.2 Unseen Data
Deep learning-based FDD algorithms can not perform well on unseen operation modes and faults. To deal with previously unknown faults, an algorithm based on supervised contrastive learning (SCL) is used in [64], in which the normal samples are contrasted with negative augmentations of themselves. Contrastive Learning is a deep learning technique for unsupervised representation learning. The goal is to learn a representation of data such that similar instances are close together in the representation space, while dissimilar instances are far apart. Soft Brownian Offset (SBO), sampling and shrinkage autoencoder are used to generate the out-of-distribution (OOD) samples, to obtain some information from unknown classes. The goal is to prevent unseen fault samples from being classified as normal conditions. Balanced fault diagnosis, imbalanced fault diagnosis, and few-shot diagnosis of TEP are tested by the proposed method, demonstrating good performances. Faults 1, 2, 4, 6, and 7 are considered as the seen faults, and faults 8, 12, 14, and 18 are selected as the unseen faults. The detection accuracies in the balanced scenario are 97.5%, 100%, 100%, and 90.63%. In the imbalanced scenario, where the balance ratio of normal to fault is 10:1, the accuracies are 97.2%, 98.75%, 100%, 90.38%, In few-shot learning, where there is only 15 samples of each fault, the accuracies are 95.1%, 97.6%, 99.86%, 88.2% respectively. Xiao et al. [94] implement a developed domain generalization (DG) method that guarantees generalization over unseen domains. DG learns a universal fault diagnosis model from historical operating modes and can generalize to unseen modes. A novel labelling and class progressive adversarial learning (LCPAL) algorithm for FDD is introduced to learn domain-invariant feature representations. This model aligns the label distribution by weighting the source classification loss based on the number of instances in each category. Stacked autoencoders can be used for handling incomplete, missing or unlabeled data. In [25] a modified stacked autoencoder is used to learn low-feature representations from incomplete data input. Only 40 input variables out of 52 are randomly chosen for fault detection and the average accuracy for the fault diagnosis of faults 1, 4, 7, 8, 10, 13 is 83.76%.
5.3 Imbalanced Data
Industrial systems usually work in normal condition, and normal data is ample. However, faults occur infrequently and the number of fault samples is rarely equal to the number of normal samples. AI algorithms can not perform well when the data is imbalanced. The performance of the data-driven FDD algorithms depends on the availability of the data, however, in many real-world cases, data is imbalanced or incomplete. To deal with this problem and achieve good FDD accuracy, different approaches are introduced. One approach for handling imbalanced data sets is generating synthetic samples of the minor class to make the data balanced [3]. The other approach is using feature extraction methods focusing on available data and considering one-class classification and anomaly detection methods.
Tian et al.[82], used a generative adversarial network (GAN) to reconstruct missing data, Spearman’s rank correlation coefficient (SRCC) for dimension reduction, select the most important features and noise cancelling, and finally DBN for FDD. The average FDR of the process for 20 faults of TEP is 89.70% which is 6.23% higher than the traditional DBN method. Hu et al.[30], proposed an incremental imbalance modified deep neural network (incremental-IMDNN) in combination with active learning for FDD of imbalanced data. An imbalance modified synthetic minority oversampling technique (IM-SMOTE) is used for addressing data imbalance. The model extends the structure hierarchically with the arrival of new faults. Similar faults are merged into the same classes by using fuzzy clustering. To tackle the class imbalance problem in FDD of chemical processes, an Imbalance Modified Convolutional Neural Network combined with incremental learning, as described in [24], generates new samples from imbalanced raw data using a dynamic modified synthetic minority oversampling technique (DM-SMOTE). This model updates itself by arriving the new data.
To identify unknown faults a few-shot learning-based unknown recognition and classification (FSLB-UR&C) approach is introduced in [55]. For feature extraction and comparing the similarity of features, a gated recurrent unit with a dot-product attention layer is used. This method does not require a large amount of data for training. The classification accuracy of the known faults of TEP is 88.9%, and for unknown faults is 85.7%. The known faults include Faults 0, 2, 12, and 14, while the unkown faults include Faults 1, 4, 5, 7, 8, 10, 11, and 13. Xu et al. [97], propose a few-shot learning fault diagnosis method based on a class-rebalance strategy, to diagnose new scarce faults. The fault samples are transferred into a feature space, where each type of fault, even each new instance, is mapped into a separate feature cluster. In this way, fault diagnosis can be achieved by estimating feature similarity between instances and faults. 1D-CNN with an attention module is used for feature extraction. Data augmentation and class rebalance are done to increase the fault samples of scarce classes. Considering 3 to 50 samples of faults for each class the average diagnosis accuracy of the proposed model on TEP is from 81.8% to 94.7%.
Zhang et al. [123]proposed a zero-shot learning (ZSL) approach to handle unseen faults. Zero-shot learning involves the utilization of expert-provided high-level descriptions instead of relying on trained objects to identify target items. These descriptions encompass semantic attributes such as colour, shape, and behaviours, which can be learned beforehand without requiring samples from unfamiliar categories. In essence, ZSL is an approach to train models to recognize patterns in new types of data using information gleaned from familiar classes along with relevant descriptions. At first feature extraction from raw data is done using a one-dimensional convolutional neural network (1D CNN), and then labelling is done by semantic descriptions by expert knowledge. Five Different subsets of TEP are used to validate the proposed method, considering different faults as unseen data. The detection accuracies for different scenarios vary from 59.72% to 96.67%, which outperforms other ZSL methods like SAE in most of the scenarios.
To address the class imbalance problem double branch rebalanced network (DBRN) method is proposed in [52]. At first, a resampling method rebalances the classes, then a cost adaptive reweighting strategy rebalances the cost of each branch. Finally the two aforementioned steps are merged by a fusing learning strategy. The imbalance ratio is 10:1, and for faults 1,2,6 of TEP the average diagnosis accuracy is 98.75%.
Guo et al. [25], used a modified stacked autoencoder to learn low-level features containing the information of complete data based on incomplete input data. This method creates data representations that can be used for classification. The steps of the data driven fault detection process are shown in 11. When the data is not rich, two general approaches can be handle the problem. Augmenting or generating more data, or using the models that can perform with few samples. Different scenarios and possible solutions are shown in the mentioned figure. Table 5 shows the summary of methods that can handle imbalanced, incomplete or unlabeled data.
Method |
Reference |
Citation |
Specification |
Fault Information |
Average Accuracy |
Unlabelled |
[57] |
Morales 2019 |
Combination of autoencoder and LSTM model and a control chart for anomaly detection of unlabelled data |
6.25% of the data is labelled |
97.46% |
Unlabelled |
[67] |
Peng 2020 |
Bidirectional gated recurrent unit (BGRU), weighting training example to overcome the class imbalance and cost-sensitive active learning to explore the unlabelled data |
Data imbalance ratio of 0.5 and the unlabelled proportion data of 0.6 |
97.50% |
Unlabelled and Imbalanced |
[65] |
Peng 2020 |
Enhanced auxiliary classifier generative adversarial network (EACGAN) to deal with imbalanced data and more stable training |
Normal to fault ratio of 10:1, all faults except 3, 9 and 15 |
77.01% |
Imbalanced |
[128] |
Zheng 2022 |
Three-step high-fidelity positive-unlabeled (THPU) is used to deal with having a few labels in the data set, by generating and augmenting data |
One labelled normal sample is available the rest is unlabelled |
92.32% |
Unlabelled and Imbalanced |
[63] |
Peng 2022 |
Contrastive learning for information extraction of unlabeled and imbalanced data |
Only 20 samples of each fault is available |
86.59% |
Imbalanced |
[4] |
Amini 2022 |
Source-aware autoencoder (SAAE) for handling imbalanced data combining BiLSTM and ResNet |
Detecting unseen faults |
93.80% |
Unlabelled |
[51] |
Ma 2022 |
Consistency regularization autoencoder (CRAE) for handling unlabelled data |
All faults except 3, 9, and 15 (20% of data is labelled) |
92.95% |
Unlabelled and Imbalanced |
[66] |
Peng 2022 |
Using genetic algorithm and sparse autoencoder for imbalanced and unlabeled fault diagnosis |
Unlabelled data and the normal to fault ratio of 10:1 |
92.02% |
Imbalanced |
[72] |
Qin 2022 |
High-efficiency GAN using Wasserstein GAN and Auxiliary Classifier, to generate and classify data for imbalanced dataset |
Normal to fault ratio of 5:1 |
82.00% |
Unlabelled |
[122] |
Zhang 2022 |
Combined LSTM for temporal feature extraction and ladder autoencoder (LAE) to handle unlabeled data |
All faults except 3,13, 15 and 16, 11% of the dataset is labelled |
95.90% |
Unknown Faults |
[55] |
Mirzaei 2023 |
Few-shot learning and GRU with attention layer for unknown fault identification |
Faults 0, 2, 12, and 14 as known faults, Faults 1, 4, 5, 7, 8, 10, 11, and 13 as unknown |
85.70% |
Imbalanced |
[97] |
Xu 2023 |
Few-shot learning based on a class-rebalance strategy to detect new faults. Feature extraction is done using 1D-CNN with an attention module |
3 to 50 samples of faults for each class |
From 81.8% to 94.7% |
Imbalanced |
[52] |
Ma 2023 |
Double branch rebalanced network (DBRN) to deal with imbalance classes using resampling and reweighting |
Imbalance ratio is 10:1 for faults 1,2,6 |
98.75% |
5.4 Ensemble Learning
Ensemble learning is a technique that combines the predictions of multiple models to improve the overall performance and robustness of a predictive model. The idea behind ensemble learning is that by aggregating the outputs of several models, you can often achieve better results than any single model can produce [74]. Ensemble methods work by creating a committee of base models, which can be of the same or different types, and then combining their outputs in various ways. Ensemble learning improves the accuracy, robustness and stability of the model. Plakias et al. [69], introduced an unsupervised fault detection approach by the ensemble of different autoencoders, each trained for detecting one type of fault. The final decision is made by a soft voting process. In this approach an anomaly indicator is generated for unknown samples, showing the confidence of the network regarding the condition of the system. To improve the feature learning for complex, nonlinear and high-dimensional signals, a hybrid model of integration of 1-DCNN and stacked denoising auto-encoders (SDAE) is introduced in [115]. This method has higher classification accuracy compared to 1-DCNN, SDAE and DBN. Typically, multiple ML algorithms are combined in ensemble learning; However, as a future direction of research for complex systems like TEP, ensemble learning could be used to aggregate model-based and signal-based methods. This approach could increase the robustness of the condition monitoring systems against uncertainties like imbalanced, incomplete or unlabelled data.
5.5 Explainability
AI models have become crucial tools for condition monitoring of industrial processes, providing unprecedented advancements in efficiency optimization, prediction performance, and data-driven decision making [99]. However, the integration of AI has not been without drawbacks. A significant concern arises from the inherent complexity of AI models, often treated as black boxes due to their inexplicable internal mechanisms. The inability to decipher the rationale behind AI predictions and decisions has led to a lack of trust, hindering their effective real-world implementation. In industrial contexts, where decisions have tangible consequences, the absence of interpretability becomes a critical barrier [131]. Explainable AI (XAI) methods address the interpretability issues of AI and other black-box models by providing end-users with transparent and understandable information regarding the model’s internal decision mechanisms.
XAI is an area of research focused on clarifying the intricate functionality of AI models and illuminating their decision-making processes. In essence, XAI bridges the gap between complex AI methodologies and human comprehensibility. By examining a model’s outputs and unravelling its internal workings, XAI techniques empower users with insights into the rationale behind AI-generated outcomes. Many XAI techniques have been explored in the literature, though they can generally be classified as either intrinsic approaches, where the model is designed or refined to be particularly interpretable by humans, or post-hoc approaches, where external tools are applied in parallel to analyze the model’s internal mechanisms with respect to predicted outputs [59]. Intrinsic methods, sometimes referred to as interpretable AI, are concerned with the development of the AI model itself, limiting the design to simple methods such as linear regression and logic based rules. Post-hoc methods on the other hand, can be model specific or general to any Blackbox model with varying levels of explanations either locally or globally [16]. Commonly used post-hoc XAI tools include local interpretable model-agnostic explanations (LIME) [73], which works by learning local interpretable models around the prediction, and Shapley additive explanations (SHAP) [48], which uses ideas from game theory to explain model contributions.
In the study by Bhakte et al. [7], an XAI-based condition monitoring framework is introduced specifically for understanding the predictions made by deep learning-based FDD models. The application of this research provides two main benefits: it enhances user confidence by providing explanations for correct fault identifications and it empowers model developers with insights when the model identifies faults incorrectly. Their methodology, based on the Shapley value framework, uses integrated gradients to identify the variables influencing DNN predictions, which are often related to the root causes of faults. A similar study [68] presents a comprehensive two-step framework for the detection and diagnosis of faults in industrial processes, with an emphasis on providing better understandability. The first step leverages a denoising sparse autoencoder to acquire a robust and adaptable feature embedding, while the subsequent phase employs smooth integrated gradients to unveil the significance of features tied to detected faults. Additionally, [2] presents an innovative deep learning-based FDD method that harnesses XAI to refine input features based on their relevance. By utilizing the layerwise relevance propagation (LRP) algorithm, this approach not only eliminates irrelevant input variables but also offers potential insights into the root causes of detected faults. Collectively, these studies underscore the profound impact of XAI in enhancing the transparency, dependability, and interpretability of condition monitoring models, ultimately yielding more effective fault detection, diagnosis, and comprehension.
5.6 Causality
Causality plays a key role in the condition monitoring of industrial processes, offering a better understanding of the true cause-and-effect relationships within complex systems. While XAI offers insights into the reasoning and inherent behaviour of intelligent models and their outputs, it neglects to consider the true physical systems that generate the data. In contrast, causality is an area of research that specifically focuses on understanding the true causal mechanisms driving physical systems rather than simply relying on statistical associations between variables [62]. These underlying relationships hold particular significance for industrial processes. Industrial processes are often characterized by a multitude of process variables and parameters, coupled with external disturbances and noise. Consequently, data-driven methods are highly susceptible to spurious correlations and erroneous associations. Thus, the exploration of causal analysis becomes critical for analyzing intricate industrial processes.
Causal modelling is the formal representation of these cause-and-effect relationships, often expressed through Bayesian networks or structural equation models. These models are generally crafted by domain experts using their expertise and leveraging their intimate knowledge of the systems. However, when domain knowledge is unavailable, causal models can be approximated via randomized control trials and experimental interventions. Even so, analyzing industrial processes in these ways might prove unfeasible or unsafe. To circumvent this problem, causal discovery methods, such as those examined by Vukovic et al. [84], become indispensable. These methods employ statistical algorithms to analyze historical observational data and infer causal relationships among the complex web of variables.
The integration of causality into data-driven AI referred to as Causal AI (CAI), can provide more robust and interpretable model predictions. Since causality represents the fundamental data generation mechanisms of a system, CAI models are less biased toward training data and are more effective in new situations by applying causal generalizations from past observations [17]. Causality-informed AI methods can provide more actionable insights by pinpointing the actual causes of anomalies or issues, facilitating targeted and effective responses. Additionally, CAI approaches can reduce false positives/negatives by distinguishing between spurious correlations and true causal links. Furthermore, these methods can be easily adapted for root-cause analysis, an important aspect of fault diagnosis and condition monitoring.
In the context of industrial processes, where the consequences of incorrect diagnoses can be significant, causality becomes a cornerstone. CAI enables accurate identification of root causes and provides a clear understanding of the propagation of faults, making it an indispensable tool for robust condition monitoring.
Several notable studies demonstrate the practical application of CAI in industrial process condition monitoring. For instance, Ma et al. [50] have developed a KPI-oriented hierarchical monitoring framework that utilizes Bayesian fusion and a Granger causality method to enhance fault detection and illuminate propagation paths. Similarly, Wu et al. [91] propose a multimodal monitoring framework for fault detection and propagation analysis using a deep local adaptive network (DLAN) alongside a dedicated causal Bayesian network. Li et al. [40] employ a spatial sequence-based deep learning approach, leveraging LSTM models with attention mechanisms to predict faults and identify causal variables with remarkable accuracy. Arunthavanathan et al. [5] present an autonomous framework for online monitoring that continuously updates itself, integrating causality through variable contributions extracted by a permutation algorithm. Meanwhile, Bi et al. [9] introduce a novel causal discovery method centred on the causality-gated time-series transformer (CGTST) model, explicitly identifying causal variables and refining relationships through permutation feature importance (PFI). In contrast, Wang et al. [87] harness causality with their temporal registration network (TRN), demonstrating improved root cause localization and propagation path identification. These select studies demonstrate the potential of integrating causality to advance the field of industrial process condition monitoring, offering comprehensive insights for more effective fault detection and diagnosis.
6 Conclusions
This paper offers an insightful examination of the utilization of deep learning and machine learning algorithms for industrial plant condition monitoring. Its primary focus centres on deep learning-based methods, renowned for their exceptional accuracy and resilience in this context. The Tennessee Eastman Process serves as the principal benchmark for evaluation, and the conclusions that follow apply to similar facilities in the chemical industry as well as complex systems that are facing inherent uncertainty. This paper emphasises how crucial deep learning algorithms are to making the complex process of feature extraction simpler. These algorithms may automatically identify and extract critical features without the need for manual feature engineering or prior experience by utilising state-of-the-art methods such as autoencoders. Furthermore, this study recognises and tackles the significant difficulties associated with managing unlabeled and imbalanced data, which are widespread problems in actual industrial monitoring situations. In addition to identifying these difficulties, it offers insightful information on how to navigate and overcome them, strengthening the usefulness and resilience of deep learning-based condition monitoring in industrial settings. To enhance the interpretability of the AI systems and make the decisions made by ML algorithms more trusted, Explainable AI and Causal AI methods are explored. In essence, this research advances industrial condition monitoring by clarifying the benefits of deep learning methods, highlighting their ability to simplify complicated problems, and providing workable answers to industry problems. The results showcased here facilitate industrial plant condition monitoring, benefiting both researchers and practitioners.
Declaration of interest
The authors declare that they have no competing financial interests that could have influenced the work reported in this paper.
Acknowledgement
We would like to acknowledge the financial support of FortisBC. and Natural Sciences and Engineering Research Council (NSERC) Canada under the Alliance Grant ALLRP 557088 – 20 in this research.